Precision Builders Construction Company was incorporated by Chris Stoschiek. The following activities occurred during the year a. Received from three investors $51,000 cash and land valued at $26,000, each investor was issued 2,500 shares of common stock with a par value of $0.10 per share c. Lent $3,500 to one of the investors who signed a note due in six months. e Paid $18,000 on the note for the construction equipment in (b) lignore interest) for use in the business at a cost of $46,000, one-fourth was paid in cash and the company signed a note for the balance (due in si Required 1. Create T-accounts for the following accounts: Cash, Notes Rece record the efflects of the transaction in the appropriate T-accounts, Include good referencing for each T-account Cash Notes Receivable Beg Bal Beg Bal End Bal End Bal Equipment Land Beg Bal Beg Ba
Cash Notes Receivable Beg Bal. Beg Ba End. Bal End. Bal ent Land Beg Bal Beg Bal End Bal End Bal Notes Payable Common Stock Beg Bal Beg Bal End Bal End Bal Additional Paid-in Copital Beg Ba End Bal
End. Bal 2. Using the balances in the T-accounts, fill in the following amounts for the accounting equation: ets Liabilities Stockholders Equity 4. Compute the market value per share of the stock. (Round your answer to 2 decimal places.) Market value per share per share
Answer #1
1) T-Accounts:
DEBIT AMOUNT $ CREDIT AMOUNT $
Cash A/C
a 51000 b 11500
c 3500
d 6200
e 18000
CB 11800
Equipment A/C
b 46000
Notes Payable A/C
e 18000 b 34500
CB 44500 d 28000
Additional Paid in capital A/C
d (withdrawal) 34200 a 76750 (being for personal purpose)
CB 42550
Notes Receivable A/C
c 3500
Land A/C
a 26000
Common Stock A/C
a 250
2) The Accounting Equation:
Assets 87300 (ie. 11800+46000+3500+26000) = Liabilities 44500 + Stockholder's Equity 42800(ie. 250+42550)
3) Market Value per share = Book Value / Number of shares = (250 + 42550)/2500 = 42800/2500 = $17.12 per share